OpenStudy (anonymous):

0.5x2 = 2x + 0.5 can you please help me

OpenStudy (anonymous):

thank a lot luis

Directrix (directrix):

0.5x2 = 2x + 0.5 .5 = 1/2 So, (1/2) x^2 = 2x + 1/2 Multiply through by 2 (2) (1/2) x^2 = 2(2x) + 2(1/2) x^2 = 4x + 1 x^2 - 4x - 1 = 0 So....

OpenStudy (anonymous):

soo

Directrix (directrix):

Soooo, use the QF (quadratic formula) to compute the roots. No factoring here. x = 2 + √ 3 and x = 2 - √ 3

Directrix (directrix):

Soooo :)

OpenStudy (anonymous):

sooo thank u againn :)